To test whether a greater proportion of Canadians than Americans believe Edward Snowden is a hero, a total of 1,519 Canadians and 1010 Americans were randomly selected. Out of the total number of Canadians chosen 1018 responded he is a hero against 515 Americans who claimed he is a hero.
Calculate the difference in the sample proportion?

Select one:

a. 0.67

b. 0.51

c. 0.16

d. 0.23
Answer #1

Correct option:

c.     0.16

Explanation:

n1 = 1519

\hat{p}1 =1018/1519 = 0.67

n2 = 1010

\hat{p}2 = 515/1010 = 0.51

So,

\hat{p}1-\hat{p}2=0.67-0.51=0.16